OpenStudy (anonymous):

Select the equation of the line in slope-intercept form 4x + 9y = 36

OpenStudy (anonymous):

m=−4/9 that's the slope

OpenStudy (anonymous):

step 1: 4x+9y=36 step 2: 9y=-4x+36 - we try getting y by its self. move the 4x over the = sign. it turned into a -4x. step 3: y=-4/9x+4 - we divided everything by 9. so -4/9 is your slope

OpenStudy (anonymous):

thank you that help

OpenStudy (anonymous):

your welcome................if you need with anything math i'll love to help